OpenStudy (anonymous):

The length of a rectangle is 8 centimeters less than twice its width. The perimeter of the rectangle is 50 cm. What are the dimensions of the rectangle?

OpenStudy (anonymous):

The first thing we should ask ourselves is "What is the formula for the perimeter of a rectangle?" You may recall that this is 2L + 2W. So 2L + 2W = 50. We also know that "The length of a rectangle is 8 centimeters less than twice its width." So we need to show this algebraically. L = 2W - 8. Now, we can substitute this into our original equation, so we're left with only one variable. 2L + 2W = 50. 2(2W - 8) + 2W = 50. Now it's just a matter of simplifying our equation. 4W - 16 + 2W = 50. 6W - 16 = 50. 6W = 66. W = 11. Now we just need to figure out length. There are two ways to do this. We can use our original perimeter formula: 2L + 2W = 50. 2L + 22 = 50. 2L = 28. L = 14. Or we can use our length formula: L = 2W - 8. L = 22 - 8. L = 14. In either case, we have our answer: The dimensions of the rectangle are 14x11, or 11x14.

OpenStudy (anonymous):

Thanks!